the perimeter of a rectangle is 46cm the length of the rectangle is 4cm less than twice the width. find the width of the rectangle
Answer by macston(5194)

L=length=2W-4; W=width; P=perimeter=2(L+W)
P=2(L+W) Substitute for L, enter given value for P
46 cm=2(2W-4+W)
46 cm=2(3W-4)
46 cm=6W-8 Add 8 to each side
54 cm=6W Divide each side by 6
9 cm=W
ANSWER Width of the rectangle is 9 cm.
CHECK
L=2W-4=18-4=14 cm
P=2(L+W)
46 cm=2(14 cm+ 9 cm)
46 cm=2(23 cm)
46 cm=46 cm